OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find 2f(1) + 3g(4). 45 43 42

OpenStudy (anonymous):

are those the answer choices ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I know if you are wrong if it is not one of those three

OpenStudy (anonymous):

Ok thats very simple we know that: \[f(x) = 2x + 1\] And, \[g(x) = 3x + 1\] Therefore, \[2f(1) +3g(4) = 2(3) +3(13) = 6 + 39 = 4\] @gazed: 45 is the right answer

OpenStudy (anonymous):

okay thank you that helps me alot :)

OpenStudy (anonymous):

Typo err: \[2f(1) = 3g(4) = 2(3) +3(13) = 6 + 39 =45\]